OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation 4(2x − 3) = 20 is given below. Choose the step that is missing from the solution. 4(2x − 3) = 20 8x = 32 x = 4 6x − 3 = 24 8x − 12 = 20 2x − 12 = 24 6x + 12 = 5

OpenStudy (anonymous):

4(2x − 3) = 20 4*2x-4*3=20 8x-12=20 8x-12+12=20+12 8x = 32 8x/8=32/8 x = 4 means 8x − 12 = 20 @ahoward79 :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

ur welcome ;)